1. Explain the term Price Elasticity of Demand. When a firm raises the price of a commodity from $10 to $20, the quantity demanded falls from 10 units to 2. Calculate the price elasticity of demand using the average price method.
Answer #1
Original price = 10
New price = 20
Change in price = 20-10 = 10
Average price = 20+10/ 2 = 15
% change in price = 10/15 = 66.67%
Original demand = 10
New demand   = 2
Channge in demand = 2-10 = -8
Average demand = 10+2 /2 = 6
% change in demand = -8/6 =   -133.33%
Price elasticiity = % change in demand / % change in price
(-133.33% / 66.67%) = -2
Answer is   -2